OpenStudy (anonymous):

Domains of Rational Expressions - Solve 37 __________ k2 – k – 30

OpenStudy (anonymous):

K^2 - k - 30 = 0 is the next part, can't figure it out from there

OpenStudy (anonymous):

try factoring

OpenStudy (anonymous):

\[k^2-k-30=(k-6)(k+5)\]

OpenStudy (anonymous):

it is not the difference of two squares in this problem, but it still factors